OpenStudy (anonymous):

is X^2+7 in C[x] irreducible?

OpenStudy (amistre64):

in C(x)? no

terenzreignz (terenzreignz):

That was rather abrupt...

OpenStudy (anonymous):

can you tell me why

OpenStudy (amistre64):

(a-bi)^2 = a^2 + b^2

OpenStudy (amistre64):

lol .... i knew it was something with abis

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

i^2 = -1

terenzreignz (terenzreignz):

hey, @urbanderivative A counter-question... a polynomial p(x) is not irreducible in F[x] if F contains some roots of p(x), right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, what are the roots of x^2 + 7 ? There are at most 2.

OpenStudy (anonymous):

haha you removed the equation....

terenzreignz (terenzreignz):

Oh Cr*p... Sorry \[\large x^2+7=x^2-(i\sqrt7)^2\]

terenzreignz (terenzreignz):

Do apply difference of two squares again ;)

OpenStudy (anonymous):

haha (x-sqr(7)i)(x+sqr(7)i)

terenzreignz (terenzreignz):

That's right :) So it can be factored into... \[\huge (x+i\sqrt7)(x-i\sqrt7)\]So... it can be factored into polynomials... the question is, are these polynomials both in \[\huge \mathbb{C}[x]\]?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Therefore, is x^2 + 7 irreducible in C[x] ? I didn't think so ;)

OpenStudy (anonymous):

haha yes, thank you

terenzreignz (terenzreignz):

By the way, all polynomials with real coefficients have roots in C. So, the only polynomials with real coefficients that are irreducible in C are linear or constants.

terenzreignz (terenzreignz):

^Follows from the fundamental theorem of Algebra :) Have fun :D